OpenStudy (krispy):

Solve for p. 3+|1+p|=12 A.p = 8 or p=−10 B.p = 14 or p=−16 C.p = 10 or p=−10 D.p = 8 or p=−8

OpenStudy (venomblast):

subtract the 3 and you will be left with |1+p|=9 so the rules with absolute value you remove the bar and solve for me and for the other equation you set the answer negative 1+p=9 or 1+p=-9

OpenStudy (venomblast):

Answer choice c